OpenStudy (anonymous):

Express using positive exponents. Then, if possible, simplify (−3)^−4

OpenStudy (anonymous):

-7

OpenStudy (anonymous):

subtract 4 do you understand?

OpenStudy (anonymous):

(-3)^-4

OpenStudy (anonymous):

do it

OpenStudy (anonymous):

yes it is exponents -4

OpenStudy (anonymous):

solve it

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

do you know how to work it out

OpenStudy (anonymous):

you are not being helpful at all @ilikephysics2

OpenStudy (anonymous):

im making him do the work they are suppose to try it, unlike you who just give them answers

OpenStudy (anonymous):

lol do you know me? im not sure how saying "do it" and solve it help ive already messaged him on how to do it

OpenStudy (anonymous):

oki need help i dont care who helps me just show me what to do

OpenStudy (unklerhaukus):

\[(-3)^{-4}=\frac{1}{(-3)^4}=\frac{1}{-3\times-3\times-3\times-3}\]

OpenStudy (anonymous):

thank you